OpenStudy (jiteshmeghwal9):

Plz help:) Find the value of:-\[\LARGE{[(\sqrt[n]{x^2})^{n-2}]^2.}\]

OpenStudy (jiteshmeghwal9):

I m getting \[x^2\]

OpenStudy (jiteshmeghwal9):

but in my book answer is given '1'.

OpenStudy (jiteshmeghwal9):

so I m confused:(

OpenStudy (unklerhaukus):

\[{[(\sqrt[n]{x^2})^{n-2}]^2.}\]\[=\left(x^{\frac2n}\right)^{2n-4}\]

OpenStudy (jiteshmeghwal9):

sorry the question is \[\LARGE{[(\sqrt[n]{x^2})^{n/2}]^2.}\]

OpenStudy (jiteshmeghwal9):

very sorry

OpenStudy (unklerhaukus):

oh

mathslover (mathslover):

u r right @jiteshmeghwal9

mathslover (mathslover):

\[\huge{(x^2)^{{\frac{1}{n}}*\frac{n}{2}*2}}\]

OpenStudy (jiteshmeghwal9):

k! thanx @mathslover @mathslover @mathslover

OpenStudy (unklerhaukus):

\[ \left[(\sqrt[n]{x^2})^{n/2}\right]^2\]\[=\left[\left(x^{\frac2n}\right)^n\right]\]\[=x^2\]

mathslover (mathslover):

i dont think that there is any case where 1 can be there ...

OpenStudy (jiteshmeghwal9):

so I m right. Phew:) thanx to all.